A. Calculate the pH of a buffer made from 0.28 M HNO2 and 0.23


M NO2-. The ka for HNO2 is 4.6 x 10-4.


B. What is the pH of the solution containing 0.20 M NH3 and 0.15


M NHACI?

1
Expert's answer
2022-04-29T13:51:03-0400

А. pH = pKa + log A/HA

pKa = logKa = log 4.6•10-4 = 3.34

pH = 3.34 + log (0.23/0.28) = 3.25


B. Kb (NH3) = 1.8•10-5 pKb= 4.74

pOH = pKb - log (NH4Cl/NH3)

pOH = 4.74 - log(0.2/0.15)= 4.62

pH = 14 - pOH = 14 - 4.62 = 9.38